# Taming OOD Actions for Offline Reinforcement Learning: An Advantage-Based Approach

### v3

@@ Abstract
Offline reinforcement learning (RL) aims to learn decision-making policies from fixed datasets without online interactions, providing a practical solution where online data collection is expensive or risky. However, offline RL often suffers from distribution shift, resulting in inaccurate evaluation and substantial overestimation on out-of-distribution (OOD) actions. To address this, existing approaches incorporate conservatism by indiscriminately discouraging all OOD actions, thereby hindering the agent’s ability to generalize and exploit beneficial ones. In this paper, we propose Advantage-based Diffusion Actor-Critic (ADAC), a novel method that systematically evaluates OOD actions using the batch-optimal value function. Based on this evaluation, ADAC defines an advantage function to modulate the Q-function update, enabling more precise assessment of OOD action quality. We design a custom PointMaze environment and collect datasets to visually reveal that advantage modulation can effectively identify and select superior OOD actions. Extensive experiments show that ADAC achieves state-of-the-art performance on almost all tasks in the D4RL benchmark, with particularly clear margins on the more challenging tasks. Our code is available at https://github.com/NUS-CORE/adac.
@@ 1 Introduction
Offline reinforcement learning (RL) lange2012batch ; levine2020offline focuses on learning decision-making policies solely from previously collected datasets, entirely circumventing the need for online interactions with the environment. This paradigm is particularly appealing for applications where online data collection is prohibitively expensive or poses safety concerns, thereby fueling its growing adoption kalashnikov2018scalable ; prudencio2023survey . While conceptually attractive, offline RL often suffers from distribution shift between the behavior policy and the learned policy. The policy evaluation on out-of-distribution (OOD) actions is prone to extrapolation error fujimoto2019off , which can be amplified through bootstrapping, leading to significant overestimation kumar2020conservative .
@@
To mitigate overestimation, a common strategy in offline RL is to incorporate conservatism into algorithm design. Value-based methods kumar2020conservative ; kostrikov2021offline ; lyu2022mildly achieve this by learning a pessimistic value function that reduces the estimated value of OOD actions to discourage their selection. Alternatively, policy-based methods fujimoto2021minimalist ; fujimoto2019off ; wang2022diffusion enforce conservatism by constraining the learned policy to remain close to the behavior policy, thereby avoiding querying OOD actions. Similarly, conditional sequence modeling approaches chen2021decision ; janner2022planning ; ajay2022conditional inherently induce conservative behavior by restricting the policy to imitate the behavior contained in the offline dataset. In a distinct manner, model-based approaches kidambi2020morel ; yu2020mopo ; sun2023model ensure conservatism by learning a pessimistic dynamics model where uncertainty-based penalization systematically underestimates the value of OOD actions.
@@
However, while conservatism is celebrated in offline RL, existing methods achieve it by indiscriminately discouraging all OOD actions, thereby hindering their capacity for generalization. Offline datasets, in practice, are usually characterized by sub-optimal trajectories and narrow state-action coverage. Consequently, an effective offline RL algorithm should possess the ability to stitch sub-optimal trajectories to generate the best possible trajectory supported by the dataset, and even extrapolate beyond the dataset to identify potentially beneficial actions. Such indiscriminate discouragement, however, severely impedes the agent’s ability to generalize and achieve high performance. This naturally leads to a fundamental question: can we develop a method that reliably distinguishes between undesirable and beneficial OOD actions, thereby balancing conservatism and generalization?
@@
To this end, we propose Advantage-based Diffusion Actor-Critic (ADAC), a novel method that systematically assesses the quality of OOD actions, selectively encourages beneficial ones, and discourages risky ones to maintain conservatism. Our key idea builds on the insight that the optimal value function provides a natural baseline for evaluating state desirability. Specifically, we regard an OOD action as advantageous if it can move the current state to a successor state whose value, under the optimal value function, exceeds that of any reachable state under the behavior policy. Since the true optimal value function is inaccessible from offline data, we adopt the batch-optimal value function ($V_{\mu}^{\ast}$ in Figure 1) as a principled surrogate. We show theoretically that the batch-optimal value function can be reliably approximated via expectile regression on the dataset. Based on this approximation, we define an advantage function to assess the desirability of actions, which is then used to modulate the temporal difference (TD) target during Q-function (critic) learning. The policy (actor) is parameterized as a diffusion model due to its strong expressiveness ho2020denoising ; wang2022diffusion and is updated under the guidance of the learned Q-function. We summarize the full architecture of ADAC in Figure 1. To the best of our knowledge, ADAC represents a pioneering attempt to systematically evaluate OOD actions, enabling the selection of beneficial ones. As a result, ADAC achieves state-of-the-art (SOTA) performance on the majority of the D4RL fu2020d4rl benchmark tasks.
Figure 1: Full architecture of ADAC. Training involves learning a diffusion actor and a Q-critic with advantage modulation, while inference deploys the learned diffusion policy.
@@ 2 Preliminaries
Offline Reinforcement Learning. RL problems are commonly formulated within the framework of a Markov Decision Process (MDP), defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,\rho_{0},P,\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the action space, $r(\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\to[-R_{\max},R_{\max}]$ is a bounded reward function, $\rho_{0}(\bm{s})$ specifies the initial state distribution, $P(\bm{s^{\prime}}\,|\,\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}_{+}$ defines the transition dynamics, and $\gamma\in(0,1)$ is the discount factor sutton2018reinforcement .
A policy $\pi(\cdot|\bm{s})$ maps a given state $\bm{s}$ to a probability distribution over the action space. The value function of a state $\bm{s}$ under a policy $\pi$ is the expected cumulative return when starting in $\bm{s}$ and following $\pi$ thereafter, i.e., $V^{\pi}(\bm{s})=\mathbb{E}_{\bm{a_{t}}\sim\pi(\cdot|\bm{s_{t}})}\big{[}\sum_{t=0}^{\infty}\gamma^{t}r(\bm{s_{t}},\bm{a_{t}})\,\big{|}\,\bm{s_{0}}=\bm{s}\big{]},$ where the expectation takes over the randomness of the policy $\pi$ and the transition function $P$. The optimal state value function satisfies the following Bellman optimality equation:
$V^{\ast}(\bm{s})=\max\limits_{\bm{a}\in\mathcal{A}}\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot\,|\,\bm{s},\bm{a})}\big{[}r(\bm{s},\bm{a})+\gamma V^{\ast}(\bm{s^{\prime}})\big{]}.$ (1)
@@
[…]
@@
Diffusion Model. Diffusion models sohl2015deep ; ho2020denoising ; song2020score consist of a forward process that corrupts data with noise and a reverse process that reconstructs data from noise. Specifically, the forward process is conducted by gradually adding Gaussian noise to samples $\bm{x_{0}}$ from an unknown data distribution $p_{\theta}(\bm{x_{0}})$, formulated as:
$q(\bm{x_{1:T}}\,|\,\bm{x_{0}}):=\prod_{t=1}^{T}q(\bm{x_{t}}\,|\,\bm{x_{t-1}}),\quad q(\bm{x_{t}}\,|\,\bm{x_{t-1}}):=\mathcal{N}(\bm{x_{t}};\sqrt{1-\beta_{t}}\bm{x_{t-1}},\beta_{t}\bm{I}),$ (3)
@@
where $T$ denotes the total number of diffusion steps, and $\alpha_{t}$ controls the variance of the added noise at each step $t$. The reverse process is modeled as $p_{\theta}(\bm{x_{0:T}}):=p(\bm{x_{T}})\prod_{t=1}^{T}p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}})$, and is trained by maximizing the evidence lower bound (ELBO) jordan1999introduction ; blei2017variational : $\mathbb{E}q\left[\log\frac{p_{\theta}(\bm{x_{0:T}})}{q(\bm{x_{1:T}}\mid\bm{x_{0}})}\right]$. After training, samples can be generated by first drawing $\bm{x_{T}}\sim p(\bm{x_{T}})$ and then sequentially applying the learned reverse transitions to obtain $\bm{x_{0}}$. For conditional generation tasks, the reverse process can be extended to model $p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}},c)$, where $c$ denotes the conditioning information.
@@
Expectile Regression. Expectile regression has been extensively studied in econometrics newey1987asymmetric and recently introduced in offline RL kostrikov2021offline . The $\tau$-expectile (with $\tau\in(0,1)$) of a real-valued random variable $\bm{X}$ is defined as the solution to the asymmetric least squares problem:
$\mathop{\arg\min}\limits_{y\in\mathbb{R}}\mathbb{E}_{x\sim X}\big{[}L_{2}^{\tau}(x-y)\big{]},$ (4)
@@
where $L_{2}^{\tau}(u)=|\tau-\mathds{1}(u<0)|u^{2}$. Therefore, $\tau=0.5$ corresponds to the standard mean squared error loss, while $\tau>0.5$ downweights the contributions of $x$ values smaller than $y$ and assigns greater weight to larger values. As $\tau\to 1$, the solution to Eq. (4) asymptotically approaches the maximum value within the support of $X$.
3 Theoretical Foundation for Advantage-Based Evaluation
In this section, we establish the theoretical foundation for our method.
3.1 Batch-Optimal Value Function Learning via Expectile Regression
In this subsection, we introduce an advantage function designed to evaluate all actions, including OOD actions. To this end, we first approximate the optimal value function, as it serves as a natural reference for assessing state desirability. Since the optimal value function satisfies the Bellman optimality equation defined in Eq. (1), we adopt expectile regression to approximate the maximum operator, while replacing the expectation with empirical samples drawn from the offline dataset. This naturally suggests the following regression problem:
@@ 3.1 Batch-Optimal Value Function Learning via Expectile Regression
Notably, this formulation entirely circumvents the need to query unseen actions. Let $V_{\tau}(\bm{s})$ denote the minimizer of $\mathcal{L}(V)$ defined in Eq. (5). The following proposition characterizes its analytical form.
@@
where $\mathbb{E}^{\tau}_{x\sim X}[x]$ denotes the $\tau$-expectile of a real-valued random variable $X$.
We further define $V_{\mu}^{\ast}(\bm{s})$ as the batch-optimal value function, which satisfies
$V_{\mu}^{\ast}(\bm{s}):=\max\limits_{\bm{a}\sim\mu(\cdot|\bm{s})}\bigg{[}r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\big{[}V^{\ast}_{\mu}(\bm{s^{\prime}})\big{]}\bigg{]}.$ (7)
The batch-optimal value function defined in Eq. (7) differs from the optimal value function in Eq. (1) in that it restricts the maximization to actions observed in the batch dataset. Therefore, it serves as a principled approximation to the optimal value function over the support of the batch data. The next proposition establishes the relation between the minimizer $V_{\tau}(\bm{s})$ and the batch-optimal value function $V_{\mu}^{\ast}(\bm{s})$.
@@
The solution $V_{\tau}(\bm{s})$ to Eq. (5) is uniformly bounded and monotonically non-decreasing with respect to $\tau$. Its pointwise limit as $\tau\to 1$, denoted by $\bar{V}(\bm{s})$, satisfies
@@
In the case where the transition dynamics P are deterministic, the limit coincides with the batch-optimal value function, i.e.,
$\lim_{\tau\to 1}V_{\tau}(\bm{s})=\bar{V}(\bm{s})=V_{\mu}^{\ast}(\bm{s}).$
According to Proposition 3.2, under deterministic transitions, $V_{\tau}(\bm{s})$ converges to $V_{\mu}^{\ast}(\bm{s})$ as $\tau\to 1$, thereby offering a precise approximation to $V^{\ast}(\bm{s})$ when the dataset $\mathcal{D}$ provides complete coverage over the action space. In contrast, under stochastic transitions, the limit $\bar{V}(\bm{s})$ involves a maximization over next states, which can lead to $\bar{V}(\bm{s})\geq V_{\mu}^{\ast}(\bm{s})$. Nonetheless, since $V_{\tau}(\bm{s})$ is monotonically non-decreasing in $\tau$, choosing $\tau<1$ yields a practical approximation to $V_{\mu}^{\ast}(\bm{s})$. This is further corroborated by our experiments, which show the best performance occurs for $\tau$ in $(0.6,0.8)$.
The above analysis justifies the approximation of the optimal value function via Eq. (5), which is subsequently used to assess the quality of OOD actions.
3.2 Advantage-Based Operator Analysis
As shown in the previous subsection, given a specific choice of $\tau$, we denote the learned value function from Eq. (5) as $V(\bm{s})$, which approximates the batch-optimal value function. Our idea is based on the observation that the quality of an action can be assessed by whether it transitions to a more desirable next state. Accordingly, we define the advantage of an action $\bm{a}$ (possibly OOD) over the behavior policy at state $\bm{s}$ as
$A(\bm{a}|\bm{s}):=\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot\mid\bm{s},\bm{a})}V(\bm{s^{\prime}})-\mathop{\text{Quantile}}\limits_{\kappa}\left(\left\{\mathbb{E}_{\bm{s^{\prime}_{i}}\sim P(\cdot\mid\bm{s},\bm{a_{i}})}V(\bm{s^{\prime}_{i}})\right\}_{i=1}^{N}\right),\quad\bm{a_{i}}\sim\mu(\cdot|\bm{s}),$ (9)

### v4

@@ Abstract
Offline reinforcement learning (RL) learns policies from fixed datasets without online interactions, but suffers from distribution shift, causing inaccurate evaluation and overestimation of out-of-distribution (OOD) actions. Existing methods counter this by conservatively discouraging all OOD actions, which limits generalization. We propose Advantage-based Diffusion Actor-Critic (ADAC), which evaluates OOD actions via an advantage-like function and uses it to modulate the Q-function update discriminatively. Our key insight is that the (state) value function is generally learned more reliably than the action-value function; we thus use the next-state value to indirectly assess each action. We develop a PointMaze environment to clearly visualize that advantage modulation effectively selects superior OOD actions while discouraging inferior ones. Moreover, extensive experiments on the D4RL benchmark show that ADAC achieves state-of-the-art performance, with especially strong gains on challenging tasks. Our code is available at https://github.com/NUS-CORE/adac.
@@ 1 Introduction
Offline reinforcement learning (RL) (Lange et al., 2012; Levine et al., 2020) focuses on learning decision-making policies solely from previously collected datasets, without online interactions with the environment. This paradigm is particularly appealing for applications where online data collection is prohibitively expensive or poses safety concerns (Kalashnikov et al., 2018; Prudencio et al., 2023). However, offline RL often suffers from distribution shift between the behavior policy and the learned policy. The policy evaluation on out-of-distribution (OOD) actions is prone to extrapolation error (Fujimoto et al., 2019), which can be amplified through bootstrapping, leading to significant overestimation (Kumar et al., 2020).
@@
To mitigate overestimation, a common strategy in offline RL is to incorporate conservatism into algorithm design. Value-based methods (Kumar et al., 2020; Kostrikov et al., 2021; Lyu et al., 2022) achieve this by learning a pessimistic value function that reduces the estimated value of OOD actions to discourage their selection. Alternatively, policy-based methods (Fujimoto & Gu, 2021; Fujimoto et al., 2019; Wang et al., 2022) enforce conservatism by constraining the learned policy to remain close to the behavior policy, thereby avoiding querying OOD actions. Similarly, conditional sequence modeling approaches (Chen et al., 2021; Janner et al., 2022; Ajay et al., 2022) inherently induce conservative behavior by restricting the policy to imitate the behavior contained in the offline dataset. In a distinct manner, model-based approaches (Kidambi et al., 2020; Yu et al., 2020; Sun et al., 2023) ensure conservatism by learning a pessimistic dynamics model where uncertainty-based penalization systematically underestimates the value of OOD actions.
@@
While conservatism is celebrated in offline RL, existing methods achieve it by indiscriminately discouraging all OOD actions, thereby hindering their capacity for generalization. Offline datasets, in practice, are usually characterized by sub-optimal trajectories and narrow state-action coverage. Consequently, an effective offline RL algorithm should possess the ability to stitch sub-optimal trajectories to generate the best possible trajectory supported by the dataset, and even extrapolate beyond the dataset to identify potentially beneficial actions. Such indiscriminate discouragement, however, severely impedes the agent’s ability to generalize and achieve high performance. This naturally leads to a fundamental question: How can we reliably distinguish between undesirable and beneficial OOD actions, and advance the trade-off between conservatism and generalization?
@@
To this end, we propose Advantage-based Diffusion Actor-Critic (ADAC), a novel method that more reliably assesses the quality of OOD actions, selectively encourages beneficial ones, and discourages risky ones. Our key insight is that the (state) value function is generally learned more reliably than the action-value function, given limited offline data; we thus use the next-state value to indirectly assess each action. Specifically, we regard an OOD action as advantageous if it can move the current state to a successor state whose value, under the optimal value function, exceeds that of any reachable state under the behavior policy. Since the true optimal value function is inaccessible from offline data, we adopt the dataset-optimal value function (optimal with respect to the offline dataset, see $V_{\mu}^{\ast}$ in Figure˜1 and the definition in Eq. (5)) as an approximation. We provide theoretical insights showing that the dataset-optimal value function can be reliably approximated through expectile regression on the dataset. Based on this approximation, we define an advantage function to assess the desirability of actions. It is then used to modulate the temporal difference (TD) target more discriminatively during Q-function (critic) learning. To have a fair comparison, we parameterize the policy (actor) using diffusion models (Ho et al., 2020; Wang et al., 2022) and is updated under the guidance of the learned Q-function. Overall, ADAC evaluates OOD actions more reliably than prior works and achieves state-of-the-art (SOTA) performance on the majority of the D4RL (Fu et al., 2020) benchmark tasks.
Figure 1: ADAC Architecture: We use an approximated optimal value function (learned with a batch data) as the measure to evaluate OOD actions. We use an approximate optimal value function (learned from batch data) to evaluate OOD actions. Its relative advantage over in-distribution actions is then used to modulate critic update.
@@ 2 Preliminaries
Offline Reinforcement Learning. RL problems are commonly formulated within the framework of a Markov Decision Process (MDP), defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,\rho_{0},P,\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the action space, and $r(\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\to[-R_{\max},R_{\max}]$ is a bounded reward function with $R_{\max}$ being the maximum absolute value of the reward. $\rho_{0}(\bm{s})$ specifies the initial state distribution, $P(\bm{s^{\prime}}\,|\,\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}_{+}$ defines the transition dynamics, and $\gamma\in(0,1)$ is the discount factor (Sutton & Barto, 2018).
A policy $\pi(\cdot|\bm{s})$ maps a given state $\bm{s}$ to a probability distribution over the action space. The value function of a state $\bm{s}$ under a policy $\pi$ is the expected cumulative return when starting in $\bm{s}$ and following $\pi$ thereafter, i.e., $V^{\pi}(\bm{s})=\mathbb{E}_{\bm{a_{t}}\sim\pi(\cdot|\bm{s_{t}})}\big[\sum_{t=0}^{\infty}\gamma^{t}r(\bm{s_{t}},\bm{a_{t}})\,\big|\,\bm{s_{0}}=\bm{s}\big],$ where the expectation takes over the randomness of the policy $\pi$ and the transition dynamics $P$. The optimal state value function $V^{*}(\cdot)$ satisfies the following Bellman optimality equation:
$V^{\ast}(\bm{s})=\max\limits_{\bm{a}\in\mathcal{A}}\big[r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot\,|\,\bm{s},\bm{a})}V^{\ast}(\bm{s^{\prime}})\big].$ (1)
@@
$J(\pi)=\int_{\mathcal{S}}\rho_{0}(\bm{s})V^{\pi}(\bm{s})\,ds=\mathbb{E}_{\bm{s_{0}}\sim\rho_{0},\bm{a_{t}}\sim\pi,\bm{s_{t+1}}\sim P}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\bm{s_{t}},\bm{a_{t}})\right].$ (2)
@@
Diffusion Model. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) consist of a forward process that corrupts data with noise and a reverse process that reconstructs data from noise. Specifically, the forward process is conducted by gradually adding Gaussian noise to samples $\bm{x_{0}}$ from an unknown data distribution $p_{\theta}(\bm{x_{0}})$, formulated as:
$q(\bm{x_{1:T}}\,|\,\bm{x_{0}})=\prod_{t=1}^{T}q(\bm{x_{t}}\,|\,\bm{x_{t-1}}),\quad q(\bm{x_{t}}\,|\,\bm{x_{t-1}})=\mathcal{N}(\bm{x_{t}};\sqrt{1-\beta_{t}}\bm{x_{t-1}},\beta_{t}\bm{I}),$ (3)
@@
where $T$ denotes the total number of diffusion steps, and $\beta_{t}$ controls the variance of the added noise at each step $t$. The reverse process is modeled as $p_{\theta}(\bm{x_{0:T}}):=p(\bm{x_{T}})\prod_{t=1}^{T}p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}})$, and is trained by maximizing the evidence lower bound (ELBO) (Blei et al., 2017): $\mathbb{E}_{q}\left[\log\frac{p_{\theta}(\bm{x_{0:T}})}{q(\bm{x_{1:T}}\mid\bm{x_{0}})}\right]$. After training, samples can be generated by first drawing $\bm{x_{T}}\sim p(\bm{x_{T}})$ and then sequentially applying the learned reverse transitions to obtain $\bm{x_{0}}$. For conditional generation tasks, the reverse process can be extended to model $p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}},c)$, where $c$ denotes the conditioning information.
@@
Expectile Regression. Expectile regression has been extensively studied in econometrics (Newey & Powell, 1987) and recently introduced in offline RL (Kostrikov et al., 2021). The $\tau$-expectile (with $\tau\in(0,1)$) of a real-valued random variable $x$ is defined as the solution to the asymmetric least squares problem:
$\mathop{\arg\min}\limits_{y\in\mathbb{R}}\mathbb{E}_{x}\big[L_{2}^{\tau}(x-y)\big],$ (4)
@@
where $L_{2}^{\tau}(u)=|\tau-\mathds{1}(u<0)|u^{2}$. Therefore, $\tau=0.5$ corresponds to the standard mean squared error loss, while $\tau>0.5$ downweights the contributions of $x$ values smaller than $y$ and assigns greater weight to larger values. Note that as $\tau\to 1$, the solution to Eq. (4) asymptotically approaches the maximum value within the support of $x$.
3 Theoretical Insight for Advantage-Based Evaluation
Our high-level insight is that in offline learning based on limited dataset, the (state) value function is generally learned more reliably than the action-value function, since the data of the latter is a subset of the former. Therefore, we can better evaluate each action indirectly using the next-state value. We start with approximating the optimal value function in the subsection below.
3.1 Dataset-Optimal Value Function
Since we only have limited data, we define the following dataset-optimal value function (Lyu et al., 2022):
$V_{\mu}^{\ast}(\bm{s})=\max\limits_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{a}\sim\mu(\cdot|\bm{s})}}\left[r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\left[V^{\ast}_{\mu}(\bm{s^{\prime}})\right]\right].$ (5)
It differs from the optimal value function Eq. (1) in that it restricts the maximization over actions from behavior policy, that is, the value function of the optimal policy in the dataset. In practice, it can be evaluated by maximizing over sampled data pairs in the offline dataset. Specifically, we adopt expectile regression to approximate the maximum operator, while replacing the expectation with empirical samples drawn from the offline dataset $\mathcal{D}$. Then we solve the following regression problem:
@@ 3.1 Dataset-Optimal Value Function
The minimizer of $\mathcal{L}(V)$ is characterized in the following proposition.
@@
where $\mathbb{E}^{\tau}_{x}[x]$ denotes the $\tau$-expectile of a real-valued random variable $x$.
The next proposition characterizes how $V_{\tau}(\bm{s})$ approximates the dataset-optimal value function $V_{\mu}^{\ast}(\bm{s})$.
@@
The solution $V_{\tau}(\bm{s})$ to Eq. (6) is uniformly bounded and monotonically non-decreasing with respect to $\tau$. Furthermore, $V_{\tau}(\bm{s})\to\bar{V}(\bm{s})$ pointwisely as $\tau\to 1$, where $\bar{V}(\bm{s})$ is given by
@@
In the case of deterministic transition dynamics, the limit coincides with the dataset-optimal value function Eq. (5), i.e., $\bar{V}(\bm{s})=V_{\mu}^{\ast}(\bm{s})$.
Proposition 3.2 shows that for deterministic transition dynamics, $V_{\tau}(\bm{s})$ converges to $V_{\mu}^{\ast}(\bm{s})$ as $\tau\to 1$. For generic stochastic transition dynamics, since the limit $\bar{V}(\bm{s})$ involves a maximization over next states (see Eq. (8)), it is possible that $\bar{V}(\bm{s})\geq V_{\mu}^{\ast}(\bm{s})$. In practice, given that $V_{\tau}(\bm{s})$ is monotonically non-decreasing in $\tau$, we have achieve a good approximation of $V_{\mu}^{\ast}(\bm{s})$ by choosing some $\tau<1$.
Therefore, by solving the regression problem Eq. (6), we can approximate the (dataset-)optimal value function. It is subsequently used to evaluate the quality of OOD actions indirectly.
3.2 A New Advantage Function
Our idea is based on the observation that the quality of an action can be better assessed by whether it transitions to a next state with higher value. Specifically, denote the learned value function from solving Eq. (6) by $V(\bm{s})$, and we define the advantage of an action $\bm{a}$ (possibly OOD) over the behavior policy at state $\bm{s}$ as
$A(\bm{a}|\bm{s}):=\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot\mid\bm{s},\bm{a})}V(\bm{s^{\prime}})-\text{Quantile}_{\kappa}\left(\left\{\mathbb{E}_{\bm{s^{\prime}_{i}}\sim P(\cdot\mid\bm{s},\bm{a_{i}})}V(\bm{s^{\prime}_{i}})\right\}_{i=1}^{N}\right),\quad\bm{a_{i}}\sim\mu(\cdot|\bm{s}),$ (9)